Determine the Kb of a weak base if a 0.52 M solution of the base has a pH of 10.68 at 25°C.

Kb = ? × 10 ?
(Enter your answer in scientific notation.)
Answer #1

Answer given pH = 10.68 pol = 14-10.68 = 3132 .: [OH] = 10 3:32 = 4178x169 cm ase B. who Let the us considen dissociation thegiven . c = 0.52 M a kb 2 (4.78*10-4.72 is 0.52 = 4.39 X 107 oko= 4:39 x 107 cs Scanned with CamScanner